How many liters of 70% acid solution must be mixed with 10% acid solution to get 480 L of a 60% acid solution
Answer by stanbon(75887)

How many liters of 70% acid solution must be mixed with 10% acid solution to get 480 L of a 60% acid solution
------
Using 2 variables:
Quantity Eq::: x + y = 480 L
Acid Eq::::: 70x+10y = 60*480
--------------------------------
Modify for elimination:
70x + 70y = 70*480
70x + 10y = 60*480
-------------------
Subtract and solve for "y":
60y = 10*480
-----
y = 80 L (amt. of 10% solution needed)
----
x = 400 L (amt. of 70% solution needed)
==========================================
Cheers,
Stan H.
=========